OpenStudy (shamil98):

\[\Large \int\limits_{0}^{1} \frac{ \tan^{-1} x }{ x^{2/3} } dx\] Given the integral above equals to \[\pi (\frac{ a }{ b } - \frac{ \sqrt{c} }{ d }) + \frac{ e }{ f } \ln (g)\] where a,b,c,d,e and f are positive integers with prime g and \[\gcd(a,b) = \gcd(c,d) = \gcd(e,f) = 1\] \[\text {Find the value of the 7 digit integer,} \space \frac{ }{ abcdefg }\] @kainui

OpenStudy (shadowlegendx):

This looks easy

OpenStudy (shamil98):

i have no idea how to do this one tbh

OpenStudy (kainui):

I guess I'm gonna try to bust out differentiation under the integral sign not sure how or what else to do lol.

OpenStudy (rizags):

Are a, b, c, d, e, f, and g necessarily different numbers?

OpenStudy (astrophysics):

Did you try by parts

OpenStudy (kainui):

\[\Large \int\limits_{0}^{1} \frac{ \tan^{-1} x }{ x^{2/3} } dx\] Given the integral above equals to \[\pi (\frac{ a }{ b } - \frac{ \sqrt{c} }{ d }) + \frac{ e }{ f } \ln (g)\] where a,b,c,d,e and f are positive integers with prime g and \[\gcd(a,b) = \gcd(c,d) = \gcd(e,f) = 1\] \[\text {Find the value of the 7 digit integer,} \space \frac{ }{ abcdefg }\]

OpenStudy (astrophysics):

Latex doesn't even load

OpenStudy (astrophysics):

Lol

OpenStudy (kainui):

I was just thinking the same thing

OpenStudy (shamil98):

it didn't specify whether they were diff numbers or not.

OpenStudy (rizags):

I have the answer but i evaluated the integral with a program

OpenStudy (shamil98):

i'd prefer a more human approach lol

OpenStudy (kainui):

Yeah no cheating come on lol

OpenStudy (shadowlegendx):

Multiply by cat and use BS function to take out some stuff and then bam, easy

OpenStudy (kainui):

@ShadowLegendX whot? Can we put bounds on this integral at all to help us determine the numbers? The gcd really limits it down along with g being prime and the fact that none of the digits can be 0 otherwise we wouldn't be able to determine the other number. Right now: \[0 < \pi \left( \frac{a}{b} - \frac{\sqrt{c}}{d} \right) + \frac{e}{f} \ln g \]

OpenStudy (shadowlegendx):

Kainui, kick me I deserve it.

OpenStudy (shamil98):

I was thinking of solving the integral and then setting the outcome equal to that or something.

OpenStudy (kainui):

This term right here is at most \[\frac{e}{f} \ln g \le9 \ln 7 \] I think it might be solvable without solving the integral, I mean we could do that sham but I mean it kinda defeats the point of the question you know?

OpenStudy (shamil98):

Hmm, i'm not too sure how else to go about it lol.

OpenStudy (kainui):

This gets us this inequality: \[\frac{-9 \ln 7}{\pi} \approx -7.2 \le \frac{a}{b} - \frac{\sqrt{c}}{d}\] We already know from the GCDs that we have some more restrictions on a and b along with c and d, so for instance we know a and b can't both be from the set 2,4,6,8. I mean if there's no alternate way to solve this why did they bother to go through this stupid digits garbage and why not just ask "solve the integral"? lol but maybe you're right and I have too high of expectations for this problem.

OpenStudy (shamil98):

There probably is an alternate way to solve it, some Indian guy posted it on a website for his JEE.

OpenStudy (shamil98):

I'm gonna try solving the integral see what happens lol

OpenStudy (shamil98):

Okay, so I integrated by parts and ended up with: \[\frac{ 3\pi }{ 4 } - \int\limits_{0}^{1} \frac{ 3x^{1/3} }{ 1+x^2 } dx\] the ∫ v du part ends just repeating lol

OpenStudy (irishboy123):

well that's good. find an interation....

ganeshie8 (ganeshie8):

IBP does the job here, but just for a variety you could also try the feynman way : \[ \int\limits_{0}^{1} \frac{ \tan^{-1} x }{ x^{2/3} } dx\\~\\ = \int\limits_{0}^{1}\int\limits_0^x \frac{ 1 }{(1+u^2) x^{2/3} } du\,dx\\~\\ = \int\limits_{0}^{1}\int\limits_u^1 \frac{ 1 }{(1+u^2) x^{2/3} } dx\,du\\~\\ =\int\limits_0^1 \dfrac{3(1-u^{1/3})}{1+u^2} du \] rest should be easy as you can make it a rational function by substituting \(u^{1/3}=t\)